A loan at 7% annually has an initial payment of 250, and 9 further payments. The payment amount increases by 3% each year. Find the loan balance immediately after the 7th payment.
Answer #1

A B C D E F G H I J K L WN Repayment 250 257.50 265.23 273.18 281.38 289.82 11 000 O 298.51 307.47 316.69 326.19 13 The last

Year 10 8 7 Opening Principal outstanding $304.85 $580.88 $830.24 $1,054.91 $1,256.75 $1,437.50 $1,598.77 $1,742.05 $1,868.74

Excel formula:

Yea Repayment 250 C5*(1+0.03) =C6*(1+0.03) =C7*(1+0.03) =C8*(1+0.03) =C9*(1+0.03) =C10*(1+0.03) =C11*(1+0.03) =C12*(1+0.03) =